OpenStudy (anonymous):

1 - 1/2 + 1/3 - 1/4 + 1/5 ...........upto infinity.

OpenStudy (anonymous):

@dumbcow I need a method to do this. I have one. Help?

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

alternating series, converges conditionally

OpenStudy (anonymous):

How would you prove/show it?

OpenStudy (experimentx):

do you want to calculate sum or prove it converges/ diverges ??

OpenStudy (anonymous):

Calculate its value.

OpenStudy (anonymous):

@SomeoneYouUsedToKnow : That is NOT the general term. General term : (-1)^(r-1) / r

OpenStudy (anonymous):

I have thought of one solution, need another perhaps better one.

OpenStudy (anonymous):

1 - (1/2 + 1/4 + 1/8 + ...) + (1/3 + 1/9 + 1/27+...) +(1/5 + 1/25+ ...)- (1/6 + 1/36+ ..) + ... hmm i don't think it's gonna work :/ lemme think of another one

OpenStudy (experimentx):

there's a tricky method

OpenStudy (anonymous):

Okay. Mine was the Geometrical series. @experimentX Which is?

OpenStudy (experimentx):

ln(1+x) = x/1 - x2/2 + x3/3 - x4/4 + ... + (-1)n-1.xn/n now put x=1

OpenStudy (anonymous):

Ohhhhhhh. I see.

OpenStudy (anonymous):

Thats what I got. Usint Infinite Geometrical series, Common ratio -x. Integrate from 0 to 1.

OpenStudy (anonymous):

@experimentX how do you know that result directly?

OpenStudy (experimentx):

usually do these stuffs these days.

OpenStudy (experimentx):

i would like to see your method.

OpenStudy (anonymous):

Okay. 1 - x + x^2 - x^3 ................infinity = 1 / 1+x Integrateit from 0 to 1. x - x^2 /2 +x^3 /3 - x^4 /4 ..........|0to1 = ln(1+x) |0to1 Putting the values We get req. sum inLHS and ln2 in RHS

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

ohhh nice sharan, nice

OpenStudy (anonymous):

of course, it does.

OpenStudy (anonymous):

are you preparing for the olympiads?

OpenStudy (anonymous):

Haha. IIT. you?

OpenStudy (experimentx):

another way of doing ...that's great!!!

OpenStudy (anonymous):

Haha. Thanks. I thought maybe Just maybe you could do it without any calculus.

OpenStudy (anonymous):

IIT. and IIIT.

OpenStudy (anonymous):

Try this book if you guys like solving math problems. I haven't solved it yet, but I do like it. http://blngcc.files.wordpress.com/2008/11/paulo-ney-de-souza-jorge-nuno-silva-berkeley-problems-in-mathematics.pdf

OpenStudy (anonymous):

Ahaa. Thanks. IIT and IIT Lol @ that. :D